OpenStudy (anonymous):

Indicate the equation of the line meeting the given conditions. Please put the equation in standard form. Containing A(1, 3) and B(0, 2)

OpenStudy (anonymous):

Let's use this formula\[y -y_1 = m(x-x_1)\]where \[m = {y_2 - y_1 \over x_2 - x_1}\] Then we can manipulate into standard-form, which has the form\[Ax + By = C\]